# Social learning moderates the tradeoffs between efficiency, stability, and equity in group foraging

### v1

@@ Results
In addition, we measure individual efficiency and examine its variation within a group of agents. Using the same parameter values, we plot the PDFs of the number of targets collected by each agent, as shown in Fig. 3D. From the low-$\rho$ regime to the high-$\rho$ regime, the shape of PDF becomes flatter with increasing probability of medium $N$ values, and the range of the number of targets collected by each agent narrows, indicating that the benefits from targets are distributed more equally among agents. However, while a larger $\rho$ enhances both equity and efficiency, an excessively large $\rho$ increases equity but reduces efficiency and increases burstiness.
Figure 4: Negative patches changes optimal $\rho$ and may improve efficiency (A) Efficiency as a function of density $\rho$ for three penalty coefficients (P=20P=-20, P=75P=-75, and P=200P=-200). Colors indicate the number of negative patches ($N_{-}=5$, $10$, and $15$), with the number of positive patches fixed at $N_{+}=10$. The total number of targets and negative targets in the domain is fixed at 6000. The controlling parameter for the target distribution is $\beta=2$ and the simulated group size is 50. The optimal $\rho$ increases as the penalty increases, indicating avoiding negative patches is more important than searching for targets. (B) Ignoring the penalty, scenarios containing negative patches achieve higher efficiency than those with only positive patches, indicating that adding randomly distributed fake negative patches can enhance overall efficiency.
Negative targets
@@
Social learning enables agents to aggregate and collect targets, thereby increasing efficiency. In nature, however, prey foraging for food face multifaceted risks, including predation, environmental hazards, and resource depletion, which can compromise their survival and efficiency. If agents can not only share information about the locations of beneficial resources, but also the existing risks in the domain, will the potential risks affect the results and conclusions in our collective search model with social learning? To model this, we introduce negative targets that penalize agents, while allowing agents to share information about risks, enabling others to actively avoid these negative targets after receiving social signals (See Methods). We then examine the influence of these risks on our model.
We fixed the total number of positive and negative targets at 6000 each. The number of positive patches, $N_{+}$, is set to 10. We then vary the number of negative patches, $N_{-}$, to change the ratio of positive patches to negative patches. To calculate the efficiency, we measure the number of collected positive targets ($N_{collect+}$) and negative targets ($N_{collect-}$), accounting for penalties from negative targets, using the formula $\eta=(N_{collect+}+P\cdot N_{collect-})/(N_{positive}\cdot T)$. Here, $P$ is the penalty coefficient, representing the strength of the penalty imposed by each negative target. As $P$ varies from -20 to -200, the increased penalties reduce efficiency, leading to a higher optimal $\rho$. Because larger penalized coefficients change the major concern for maximizing efficiency, shifts from rapidly collecting more positive targets to avoiding encountering negative targets, which necessitate greater $\rho$, although it will diminish the efficiency contributed from positive targets.
@@
For the same penalty coefficient $P$, we observe that the differences in efficiency curves between cases with varying numbers of negative patches increase as the penalty rises. This occurs because a higher number of negative patches allows agents to detect more negative targets, which in turn leads to a greater overall penalty.
From Fig. 4B, where the penalty for detecting negative targets is set to zero, we surprisingly find that randomly distributed negative patches significantly increase efficiency. This occurs because detected negative targets from randomly generated patches exclude certain regions of the domain, thereby constraining the area that agents explore, which reduces the time required to explore a new patch. This is true even though the patch centers are independently generated. The improvement in efficiency from adding negative patches does not appear to depend on the number of negative patches.
@@ Discussion
The presented work advances the understanding of collective foraging by demonstrating that the social learning range $\rho$—the distance over which individuals respond to social signals—acts as a crucial tuning parameter that regulates a complex interplay between efficiency ($\eta$), temporal stability, and equitable resource distribution in group foraging. Our central finding is that optimal foraging efficiency occurs at intermediate values of $\rho$, offering a unified, mechanistic explanation for patterns observed across disparate biological and computational systems [29, 27, 30, 31]. This work advances beyond previous literature by explicitly quantifying the multi-faceted costs and benefits of social information sharing in dynamic and risk-laden environments.
In network terms, $\rho$ defines the effective connectivity radius: as $\rho$ increases, each agent’s social neighborhood expands, approximating a denser communication graph. When $\rho\rightarrow 0$, the network fragments into isolated nodes; when $\rho\rightarrow 0.7$, it approaches full connectivity. Thus, the observed non-monotonic dependence of efficiency on $\rho$ parallels prior findings that intermediate network density maximizes collective performance [33, 34].
@@
Our finding of an optimal intermediate range for information sharing confirms a recurring theme in the collective search and foraging literature. The peak efficiency observed at intermediate $\rho$ aligns with the theoretical finding in Mongolian gazelles [29], where acoustic communication over intermediate distance optimizes search efficiency. This pattern of optimal intermediate communication also mirrors findings from Deborah Gordon’s work on harvester ants, where restrained signaling maintains efficient colony foraging [35, 36, 37]. Our framework generalizes these empirical observations by explicitly linking $\rho$ to the balance of three distinct behavioral modes: global exploration ($\mu=1.1$), local exploitation ($\mu=3$), and targeted walk. Furthermore, the model offer a clearer mechanistic explanation for why non-optimal range fails. When $\rho$ is too small, agents search nearly independently, leading to maximum exploration but minimum cooperation, akin to a purely individual search. This is consistent with finding that insufficient information sharing hampers successful searches [38, 39, 40]. On the other hand, if $\rho$ is too large, excessive targeted walks occur, leading to redundant search effort and premature depletion of local patches. This highlights a cost of social learning that suppresses the crucial simultaneous discovery of multiple patches [27], and effect observed in empirical studies of bird foraging [41, 42].
@@
Our model incorporates adaptive switches between two Lèvy walk modes ($\mu=1.1$ and $\mu=3$), providing a composite search strategy that advances beyond traditional Lèvy models. Pure, memoryless Lèvy walk , while optimal for sparse, randomly distributed targets [7, 2, 43], often fail to account for the adaptive behavior of real foragers. Our model supports the notion that encounter-conditional heuristics, such as Area-Restricted Search (ARS) triggered by a successful encounter, are often superior to idealized Lèvy walks, especially in patch environments [15]. Moreover, the targeted walk mode, absent from individual ARS models, serves as a socially-mediated exploitation mechanism. Our analysis of the targeted walk’s benefit-to-cost ratio provides a quantitative link between a social behavior (joining successful peers) and the resulting gain in efficiency, clearly demonstrating that the highest benefit-to-cost ratio coincides with the optimal $\rho$.
@@
Prior theoretical studies on collective foraging primarily focused on maximizing search efficiency [39, 44, 27, 30, 31]. Our work provides a missing, multifaceted perspective by explicitly quantifying the trade-offs $\rho$ creates with temporal stability and equity of resource intake. The impact of $\rho$ on the burstiness parameter and instantaneous efficiency reveals a crucial trade-off between maximizing average efficiency and maintaining a steady resource stream. While high $\rho$ can lead to short, intense bursts of instantaneous efficiency, it ultimately results in a high burstiness and increased starvation time. This bursty, highly unstable dynamics is consistent with “herding” and maladaptive informational cascades observed in human social learning experiments [45]. The lowest burstiness (maximal stability) is found at the same intermediate $\rho$ that optimizes $\eta$. This suggests that optimal efficiency is intrinsically linked to stability in dynamic foraging environments. Social learning is a “double-edged sword” , but our model shows the optimal range manages this risk, mitigating the tendency toward rapid, inflexible, and ultimately unsustainable resource exploitation that plagues systems with high social dependence [45]. The mechanism underlying the optimal intermediate $\rho$ can also be interpreted through the lens of transient diversity-the temporary coexistence of heterogeneous behaviors or information states. When $\rho$ is small, diversity persists but coordination is limited; when $\rho$ is large, information homogenizes too quickly, leading to premature convergence on local optima. Intermediate $\rho$ maintains transient diversity long enough to explore alternative solutions before consensus, echoing a general principle found across formal models of collective problem solving [46].
@@
By measuring the distribution of collected targets per agent, our work directly addresses the producer-scrounger paradigm from a spatial and kinematic perspective [27, 47]. Our results demonstrate that increasing $\rho$ flattens the distribution of collected targets among agents, thereby increasing equity. However, this gain in equity at high $\rho$ comes at the cost of overall lower efficiency and increased instability, as the collective engages in redundant search efforts. Our findings provide an individual-level mechanism for the group-size effects seen in real systems: In the Australomisidia ergandros spider, larger group sizes increase scrounger-type frequency and decrease producer frequency [47]. For human social learning, rates of copying increases with group size [45]. These experimental observations has been verified by our model. The time ratio of the targeted walk reaches 0.6 in a group of 50 agents, which is significantly higher than 0.4 in a group of 10 agents, as shown in Supplementary Figure 5, with the same target distribution.
@@
Our incorporation of negative targets (risks) into the collective search domain introduces a novel dimension of realism to foraging models. The optimal $\rho$ adpatively shift from intermediate to higher as the penalty magnitude increases. This is a key insight: when risks are severe, the group’s priority shifts from maximizing immediate resource collection to prioritizing risk avoidance through wider social information gathering. Based on that, we surprisingly found that adding randomly distributed “fake” negative patches (with zero penalty) significantly increased overall efficiency. This novel emergent mechanism arises because the negative social signals effectively constrain the search space, focusing the collective exploration effort away from empty regions, leading to faster patch discovery. This suggests a completely new principle for designing cooperative search algorithms in robotics and swarms [48, 49, 44].
@@ Methods
Variance reduction
Considering that our model is based on a random walk, which introduces randomness to the statistics, we use the same set of target distributions for cases with different $\rho$ to reduce variance. To minimize the errorbar with a 95% confidence interval, we run 2000 cases for each $\rho$.
@@
All python scripts used to generate the figures and reproduce the simulation results in this paper are openly available on our GitHub repository: https://github.com/LoneStar97/social-learning-search

### v2

@@ Results
In addition, we measure individual efficiency and examine its variation within a group of agents. Using the same parameter values, we plot the PDFs of the number of targets collected by each agent, as shown in Fig. 3D. From the low-$\rho$ regime to the high-$\rho$ regime, the shape of PDF becomes flatter with increasing probability of medium $N$ values, and the range of the number of targets collected by each agent narrows, indicating that the benefits from targets are distributed more equally among agents. However, while a larger $\rho$ enhances both equity and efficiency, an excessively large $\rho$ increases equity but reduces efficiency and increases burstiness. These results address RQ2: increasing $\rho$ simultaneously increases equity but can destabilize target intake dynamics, and the $\rho$ that maximizes mean efficiency also minimizes starvation burstiness.
Figure 4: Negative patches changes optimal $\rho$ and introducing fake negative target can improve efficiency. (A) Efficiency as a function of $\rho$ for three penalty coefficients (P=20P=-20, P=75P=-75, and P=200P=-200). Colors indicate the number of negative patches ($N_{-}=5$, $10$, and $15$), while the number of positive patches is fixed at $N_{+}=10$. The domain contains 6000 positive targets and 6000 negative targets; the target distribution is controlled by $\beta=2$, and the group size is 50. As $|P|$ increases, the efficiency-maximizing $\rho$ shifts toward larger values, indicating that broader information sharing becomes more beneficial when negative encounters carry larger penalties. (B) When the penalty is ignored ($P=0$), introducing a single negative patch ($N_{-}=1$) yields higher efficiency than the all-positive baseline ($N_{-}=0$), showing that randomly placed, uninformative negative cues can improve efficiency by constraining exploration away from unproductive regions.
Social Network Analysis
Considering that agents utilize social learning to collaborate and improve search efficiency, it is natural to use network metrics to quantify the collective dynamics of our model. To define the social network, we assume that an edge exists between two agents when their mutual distance is smaller than the social learning range $\rho$. We then extract the time-dependent network structure throughout each realization and compute the corresponding network statistics (see Supplementary Materials for details).
Among these metrics, a component is defined as a maximal set of nodes such that every pair of nodes in the set is connected by at least one path. As shown in Fig. 3E and F, when $\rho=0.01$, the PDF of the average component size per case $S$, peaks at $S=1$, while the PDF of the average number of components per case $N_{C}$, peaks at $N_{C}=50$. This indicates that agents search largely independently in most realizations. As $\rho$ increases, the PDF of $S$ shifts to the right, indicating the formation of larger connected components. Correspondingly, the number of components decreases, leading to a leftward shift of the PDF of $N_{C}$. When $\rho$ increases to $0.7$, the PDF of $S$ peaks at $S=50$ and the PDF of $N_{C}$ peaks at $N_{C}=1$, indicating that agents typically form a single, fully connected cluster.
The insets in Fig. 3E and F further show that the case-averaged component size, $\langle S\rangle$, increases monotonically with $\rho$, while the case-averaged number of components, $\langle N_{C}\rangle$, decreases monotonically with $\rho$. Notably, the parameter value corresponding to optimal search efficiency occurs at an intermediate connectivity level, with $\langle N_{C}\rangle=7.44$ and $\langle S\rangle=7.73$.
The component-based network statistics in Fig. 3E and F provide a direct network-level interpretation of the social learning range $\rho$. As $\rho$ increases, the implicit proximity-based social network transitions from a highly fragmented structure to a densely connected one. For small $\rho$, the network consists of many isolated components, indicating weak connectivity and limited information transmission between agents. As $\rho$ increases, components merge and the average component size $\langle S\rangle$ increases monotonically, while the average number of components $\langle N_{C}\rangle$ decreases. In the large-$\rho$ limit, agents typically form a single giant component, implying near-global information sharing in which information discovered by one agent rapidly propagates throughout the group. Notably, the parameter regime corresponding to maximal search efficiency coincides with an intermediate level of connectivity, where neither extreme fragmentation nor full connectivity dominates. This observation parallels results from collective decision-making models on social networks, which show that moderate connectivity often optimizes group performance by balancing information integration and independence [hahn2020truth], and contrasts with regimes of excessive connectivity that can suppress independent discovery [mann2022collective].
Beyond static connectivity, the same component statistics also reveal a dynamical mechanism consistent with the emergence of transient diversity. At low $\rho$, the persistence of many small components reflects sustained heterogeneity in agent interactions and information states; however, limited connectivity prevents efficient coordination. At high $\rho$, rapid formation of a single connected component leads to fast homogenization of information and behavior, resulting in premature convergence and redundant exploitation. In contrast, at intermediate $\rho$, the network exhibits a transiently fragmented structure: multiple components coexist during early exploration and progressively merge as information accumulates. This regime maintains heterogeneity long enough to support effective exploration while still allowing coordinated exploitation at later stages. In this sense, $\rho$ regulates not only the degree of network connectivity but also the timescale over which diversity persists. The observed efficiency maximum therefore coincides with a regime in which diversity is generated and maintained transiently rather than permanently, consistent with the general principle that transient diversity enhances collective problem solving by delaying premature consensus [smaldino2024maintaining]. Taken together, these results close RQ2 by showing that $\rho$ acts as a connectivity control parameter for the emergent proximity network: small $\rho$ yields fragmented interaction graphs and localized information flow, whereas larger $\rho$ rapidly produces a highly connected regime that homogenizes behavior. This connectivity transition mechanistically links the observed increase in equity with $\rho$ to changes in temporal stability, since highly connected networks synchronize exploitation and can amplify bursty intake dynamics. Having characterized how $\rho$ shapes efficiency and collective organization in the baseline setting, we next use the same framework in a post-hoc application to examine how the presence of risk in the environment affects collective search dynamics, with agents sharing information to avoid negative targets while maximizing positive target intake.
Collective foraging in risky environments
@@
Social learning enables agents to aggregate and collect targets, thereby increasing efficiency. In nature, however, prey foraging for food face multifaceted risks, including predation, environmental hazards, and resource depletion, which can compromise their survival and efficiency. Here we ask how such risks alter the functional role of social learning in collective search. To model this, we introduce negative targets that impose a penalty when encountered, while allowing agents to socially transmit hazard information so that nearby agents can actively avoid these locations after receiving social signals (see Methods). We then examine how risk reshapes the dependence of efficiency on the social learning range $\rho$.
We fixed the total number of positive and negative targets at 6000 each. The number of positive patches, $N_{+}$, is set to 10. We then vary the number of negative patches, $N_{-}$, to tune the spatial structure and coverage of hazardous regions. To measure the efficiency, we measure the number of collected positive targets ($N_{collect+}$) and encountered negative targets ($N_{collect-}$), accounting for penalties from negative targets, using the formula $\eta=(N_{collect+}+P\cdot N_{collect-})/(N_{positive}\cdot T)$. Here, $P$ is the penalty coefficient, representing the penalty incurred per negative target encountered. As $|P|$ increases, the efficiency-maximizing $\rho$ shifts toward larger values because socially-shared avoidance information becomes more valuable when negative encounters carry larger penalties. This favors larger $\rho$, which spreads avoidance cues more broadly, although excessively large $\rho$ can also increase redundancy in positive-target exploitation.
@@
For the same penalty coefficient $P$, we observe that the differences in efficiency curves between cases with varying numbers of negative patches increase as the penalty rises. This occurs because increasing $N_{-}$ changes the spatial coverage of hazardous regions, which increases the likelihood that agents incur penalties (or expend travel effort detouring around hazards), an effect that becomes more pronounced as $|P|$ increases.
When the penalty is ignored ($P=0$; Fig. 4B), randomly distributed negative patches can still increase efficiency. In this limit, socially shared negative cues act as exclusion signals that prune the search space, reducing exploration time before locating a new positive patch. Overall, this post-hoc application shows that when negative encounters carry penalties, broader information sharing shifts from primarily improving exploitation to actively supporting hazard avoidance, thereby pushing the efficiency-optimal $\rho$ toward larger values; when penalties are absent, “where-not-to-search” information can still improve performance by restricting exploration away from unproductive regions.
@@ Discussion
The present work advances the understanding of collective foraging by showing that the social learning range $\rho$—the distance over which individuals respond to social signals—acts as a key control parameter that jointly regulates foraging efficiency ($\eta$), temporal stability, and equity of resource acquisition. Building on the dynamical and network-level analyses above, we show that $\rho$ tunes both proximity-network connectivity and the persistence of behavioral heterogeneity through time. Our central finding is that efficiency is maximized at intermediate $\rho$, providing a mechanistic explanation for non-monotonic performance patterns reported across diverse biological and computational systems [martinez2013optimizing, bhattacharya2014collective, garg2022individual, garg2024evolution]. Beyond identifying this optimum, we explicitly quantify the coupled benefits and costs of social information sharing in environments with depleting positive targets and risky negative targets, clarifying when additional connectivity promotes coordination versus when it induces redundancy and suppresses independent discovery. From a broader network perspective, our results place range-limited social learning within a class of collective systems in which performance depends non-monotonically on effective connectivity: intermediate connectivity balances information integration with continued independent exploration, whereas overly sparse or overly dense networks degrade collective performance [hahn2020truth, mann2022collective].
@@
Our finding of an optimal intermediate range for information sharing confirms a recurring theme in the collective search and foraging literature. The peak efficiency observed at intermediate $\rho$ is consistent with the theoretical finding in Mongolian gazelles [martinez2013optimizing], where acoustic communication over intermediate distance optimizes search efficiency. This pattern of optimal intermediate communication also provides a mechanistic explanation for Deborah Gordon’s work on harvester ants, where restrained signaling maintains efficient colony foraging [gordon1996organization, gordon2013rewards, gordon2014ecology]. A similar mechanism is observed in organizational innovation networks, where excessive connectivity among inventors causes premature convergence on suboptimal solutions; placing less-imitative inventors at influential network positions helps maintain exploratory diversity and improves performance in complex landscapes [heydari2025core]. Our framework generalizes these empirical observations by explicitly linking $\rho$ to the balance of three distinct behavioral modes: global exploration ($\mu=1.1$), local exploitation ($\mu=3$), and targeted walk. Furthermore, the model offers a clearer mechanistic explanation for why non-optimal range fails. When $\rho$ is too small, agents search nearly independently, leading to maximum exploration but minimum cooperation, akin to a purely individual search. This is consistent with finding that insufficient information sharing hampers successful searches [couzin2005effective, torney2009context, rosenthal2015revealing]. On the other hand, if $\rho$ is too large, excessive targeted walks occur, leading to redundant search effort and premature depletion of local patches. This highlights a cost of social learning that suppresses the crucial simultaneous discovery of multiple patches [bhattacharya2014collective], an effect observed in empirical studies of bird foraging [ward2008quorum, biro2016bringing].
@@
Our model incorporates adaptive switches between two Lèvy walk modes ($\mu=1.1$ and $\mu=3$), providing a composite search strategy that advances beyond traditional Lèvy models. Pure, memoryless Lèvy walk , while optimal for sparse, randomly distributed targets [viswanathan1999optimizing, sims2008scaling, humphries2010environmental], often fail to account for the adaptive behavior of real foragers. Our model supports the notion that encounter-conditional heuristics, such as Area-Restricted Search (ARS) triggered by a successful encounter, are often superior to idealized Lèvy walks, especially in patch environments [ross2018general]. Moreover, the targeted walk mode, absent from individual ARS models, serves as a socially-mediated exploitation mechanism. Our analysis of the targeted walk’s benefit-to-cost ratio provides a quantitative link between a social behavior (joining successful peers) and the resulting gain in efficiency, clearly demonstrating that the highest benefit-to-cost ratio coincides with the optimal $\rho$.
@@
Prior theoretical studies on collective foraging primarily focused on maximizing search efficiency [torney2009context, torney2011signalling, bhattacharya2014collective, garg2022individual, garg2024evolution]. Our work provides a missing, multifaceted perspective by explicitly quantifying the trade-offs $\rho$ creates with temporal stability and equity of resource intake. The impact of $\rho$ on the burstiness parameter and instantaneous efficiency reveals a crucial trade-off between maximizing average efficiency and maintaining a steady resource stream. While high $\rho$ can lead to short, intense bursts of instantaneous efficiency, it ultimately results in a high burstiness and increased starvation time. This bursty, highly unstable dynamics is consistent with “herding” and maladaptive informational cascades observed in human social learning experiments [toyokawa2019social]. The lowest burstiness (maximal stability) is found at the same intermediate $\rho$ that optimizes $\eta$. This suggests that optimal efficiency is intrinsically linked to stability in dynamic foraging environments. Social learning is a “double-edged sword” , but our model shows the optimal range manages this risk, mitigating the tendency toward rapid, inflexible, and ultimately unsustainable resource exploitation that plagues systems with high social dependence [toyokawa2019social]. Taken together, our results demonstrate a concrete realization of transient diversity—the temporary coexistence of heterogeneous behaviors and information states, which emerges naturally at intermediate $\rho$. Intermediate $\rho$ maintains transient diversity long enough to explore alternative solutions before consensus, echoing a general principle found across formal models of collective problem solving [smaldino2024maintaining].
@@
By measuring the distribution of collected targets per agent, our work directly addresses the producer-scrounger paradigm from a spatial and kinematic perspective [bhattacharya2014collective, dumke2016producers]. Our results demonstrate that increasing $\rho$ flattens the distribution of collected targets among agents, thereby increasing equity. However, this gain in equity at high $\rho$ comes at the cost of overall lower efficiency and increased instability, as the collective engages in redundant search efforts. Our findings provide an individual-level mechanism for the group-size effects seen in real systems: In the Australomisidia ergandros spider, larger group sizes increase scrounger-type frequency and decrease producer frequency [dumke2016producers]. For human social learning, rates of copying increase with group size [toyokawa2019social]. These experimental observations have been verified by our model. The time ratio of the targeted walk reaches 0.6 in a group of 50 agents, which is significantly higher than 0.4 in a group of 10 agents, as shown in Supplementary Figure 5, with the same target distribution.
@@
Our incorporation of negative targets (risks) into the collective search domain introduces a novel dimension of realism to foraging models. The optimal $\rho$ adpatively shift from intermediate to higher as the penalty magnitude increases. This is a key insight: when risks are severe, the group’s priority shifts from maximizing immediate resource collection to prioritizing risk avoidance through wider social information gathering. Based on that, we surprisingly found that adding randomly distributed “fake” negative patches (with zero penalty) significantly increased overall efficiency. This novel emergent mechanism arises because the negative social signals effectively constrain the search space, focusing the collective exploration effort away from empty regions, leading to faster patch discovery. This constraint operates analogously to an informational boundary, pruning uninformative regions while preserving sufficient behavioral diversity to avoid premature convergence. This suggests a completely new principle for designing cooperative search algorithms in robotics and swarms [garnier2007biological, romanczuk2012active, torney2011signalling].
The negative-target extension should be viewed as a post-hoc application of the core framework rather than a separate research question. It illustrates how the same control parameter $\rho$ can change its functional role under additional task constraints: when hazards impose penalties, increasing $\rho$ provides value by spreading avoidance information and reducing costly encounters; when penalties are absent, negative cues can still improve efficiency by pruning unproductive regions of the search space. This framing suggests that range-limited social learning can be used diagnostically to anticipate how optimal communication scales shift when search is coupled to risk.
@@ Methods
Simulation setup and numerical implementations
Because targets are finite and non-regenerative, efficiency necessarily decreases as depletion progresses, making raw time-to-termination comparisons across conditions sensitive to the depletion level. To enable consistent comparisons across different $\rho$, we therefore evaluate performance at a standardized depletion condition and terminate each realization once agents have collected $30\%$ of the positive targets [garg2022individual, garg2024evolution]. To verify that our qualitative conclusions are not an artifact of this stopping rule, we also examine later-stage depletion by running simulations beyond the $30\%$ threshold in settings with fewer patches; these results are shown in Supplementary Fig. 4. Considering that our model is based on a random walk, which introduces randomness to the statistics, we use the same set of target distributions for cases with different $\rho$ to reduce variance. To minimize the errorbar with a 95% confidence interval, we run 2000 cases for each $\rho$.
@@
All python scripts used to generate the figures and reproduce the simulation results in this paper are openly available on our GitHub repository:
https://github.com/LoneStar97/social-learning-search.git
Additional videos of agents’ movement can be find here:
https://www.youtube.com/playlist?list=PLgRFM9nAjJRwoZvCGBAdCIE-BYNgPmSuV